Operation Summary for October 5, 2011 - October 7, 2011 |
Wednesday October 5 The day shift began with Operations monitoring the Antiproton Source (Pbar) stack (104.5E10), with the Linac, Booster, and Main Injector (MI), NuMI, and MiniBooNE all in access. During the day shift, most of the complex was in access for maintenance work and for safety system testing. NuMI personnel reported at 10:10 AM that they had repositioned their target and reconfigured their Horn power supply for antineutrino operations. At 11:08 AM, Interlock personnel reported finding a ground water leak coming from a ceiling panel in the 8 GeV transfer line. FESS reported that there were two water system leaks somewhere in the area that could be the cause: an ICW leak and a Pbar transport LCW leak (later MI personnel reported that that this leak was probably due to a Pbar magnet LCW leak). Interlocks Personnel completed their tests in the 8 GeV transfer line at 11:23 AM; FESS and MI personnel investigated the water leak. Operators rebooted TCLK for a Control person at 2:06 PM. NuMI personnel reported at 3:58 PM that the Horn target was in position for operation in the antineutrino mode. During the evening shift, by 4:43 PM, Linac, MI, and Booster were ready for beam, but Interlock personnel had to complete the Booster safety system tests. Operations established beam through Linac and Booster to MiniBooNE at 6:27 PM. Operations established beam through MI to NuMI for target scans at 7 PM. At 7:14 PM, a Linac RF station (LRF3) tripped off; a Linac person and an operator had the station back up by 7:45 PM. Operators loaded the Recycler with protons at 9:47 PM. NuMI target scans ended around 9:50 PM. Thursday October 6 The midnight shift began with Operations monitoring the Pbar stack (142.9E10) and the protons in the Recycler (37.2E10), and with MiniBooNE taking beam. During the midnight shift, operators reported tuning the 750 KeV line and the Booster at 3:30 AM to fix the drop in Booster intensity. At 7:13 AM, a Linac quadrupole magnet bulk power supply (L:Q3BULK) tripped off; it reset after a Linac technician investigated and replaced the power supply’s controls crate. During the day shift, Neutron Therapy Facility (NTF) personnel reported at 8:19 AM that they were tuning in preparation for treating one patient early in the afternoon. Operations established beam to NuMI at 9:35 AM. At 10:48 AM, MI personnel began testing a shorter ramp cycle for a new slipstacking method to improve beam efficiency. Operations established beam to the Meson FTBF experiment T-992 at noon. Booster personnel began an aperture scan study. CUB personnel reported a power glitch at 2:07 PM; they began recovery. At 3:22 PM, CUB personnel reported that column #1 was down; they will regenerate the column on Friday. The NuMI beam permit tripped off at 3:46 PM; it reset. Operators reported short downtimes due to problems with LRF3, BRF1, and BRF13. During the evening shift, Electron Cooling personnel turned off the Pelletron for the final time at 6:04 PM. At 7:05 PM, a Pbar Debuncher vacuum pump (D:IP925) tripped off; operators left it off for an hour and then reset the pump. Operations disabled beam to T-992 at 10:10 PM. Friday October 7 The midnight shift began with Operations monitoring the Pbar stack (56.4E10) and the protons in the Recycler (7.2E10), and with the NuMI and MiniBooNE experiments taking beam. During the midnight shift, operators reported at 1:01 AM that IP925 was tripping off 20 minutes after every reset. At 1:08 AM, MIF17 tripped off for the first of seven times in about two hours; it reset. Operators turned off MIRF17 at 2:51 AM and brought up MIRF15 to take its place. Operations Coordinator (OPSCO) Plans for Today and the Weekend The plans for today and the weekend are to continue delivering beam to the users. The Pbar stack will be terminate between 11 and 11:30 AM today, 10/7/11, ending the antiproton program at Fermilab. |
|
More Information Comments and Suggestions |
| last modified 10/07/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |